OpenStudy (anonymous):

MEDAL&FAN&Testimonials (n-r+1)!/(n-r-2)! How do I write this without using factorial notation?

OpenStudy (anonymous):

My head's stuck on this problem if anyone could offer me help that would be awesome!

OpenStudy (anonymous):

@ganeshie8 the almighty?

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

Do you know how to do this problem?

ganeshie8 (ganeshie8):

Hint : let \(n-r+1 = m\)

OpenStudy (anonymous):

The trickiest part is that denominator is negative whereas the numerator is positive.

OpenStudy (anonymous):

This makes me feel insecure about canceling either.

OpenStudy (anonymous):

(m)(m-1)(m-2)/(m-3)(m-4)(m-5)

ganeshie8 (ganeshie8):

\[\dfrac{(n-r+1)!}{(n-r-2)!} = \dfrac{(\color{blue}{n-r+1})!}{(\color{blue}{n-r+1}-3)!} \]

OpenStudy (anonymous):

Oh!

OpenStudy (anonymous):

Looks like the factorials are all canceled but (-3)!, at least intuitively.

ganeshie8 (ganeshie8):

Let that blue part equal \(\color{blue}{m}\) : \[\dfrac{(n-r+1)!}{(n-r-2)!} = \dfrac{(\color{blue}{n-r+1})!}{(\color{blue}{n-r+1}-3)!} =\dfrac{\color{blue}{m}!}{(\color{blue}{m}-3)!} =mP3\]

OpenStudy (anonymous):

1/-3!

OpenStudy (anonymous):

no no

OpenStudy (anonymous):

How come mP3 and not 1/-3!?

ganeshie8 (ganeshie8):

recall the definition of permutation : \[nPr = \dfrac{n!}{(n-r)!}\]

OpenStudy (anonymous):

Oh yes.

OpenStudy (anonymous):

Not the conventional calculation of fractions.

OpenStudy (anonymous):

Assuming that rule, (n+4)!/(n+2)! will be not (n+4)(n+3)?

OpenStudy (anonymous):

(n+4)(n+3)(n+2)(n+1)(n)/(n+2)(n+1)(n)=(n+4)(n+3)? This is not correct?

OpenStudy (usukidoll):

\[\frac{(n+4)(n+3)(n+2)(n+1)(n)}{(n+2)(n+1)(n)}\] cancel like terms should be (n+4)(n+3)

OpenStudy (anonymous):

At first seemed a bit counterintuitive after what Ganashie showed me.

OpenStudy (anonymous):

But I guess I am not so retarded at this timeXD